How much power would be delivered by pulling a 10.0 kg sled along a flat,snowy field at a constant speed of 1.96 m/s with a rope at an angle of 30.0°, if μk between sled and snow is 0.30?
Answer #1

here,

mass of sled ,m = 10 kg

theta = 30 degree

uk = 0.3

let the applied force be F

as the sled is moving at constant velocity

the net force on the sled is zero

F * cos(theta) - uk * ( m * g - F * sin(theta)) = 0

F * cos(30) - 0.3 * ( 10 * 9.81 - F * sin(30)) = 0

solving for F

F = 28.97 N

the power delivered , P = F * cos(theta) * v

P = 28.97 * cos(30) * 1.96 W

P = 49.2 W